The diameter of a given bare conductor is 0.50 inch. A thermoplastic insulation with thickness of 0.1 inch is wrapped around to insulate the conductor. Determine the insulation resistance of this conductor per meter. Assume specific resistance of thermoplastic insulation to be 2 x 1014 ohmcm.

1
Expert's answer
2021-12-12T18:05:11-0500

Insulation resistance R of single core cable is given by

"R=\frac{ρ}{2π l} × In\frac{r_2}{r_1}"

"\u03c1 = 2\u00d710^{14}ohm-cm = 2\u00d710^{12}ohm-m"

l = 1m

"r_1" = conductor radius = 0.5 inch =1.27cm

"r_2" = insulation radius = 0.5+0.1 = 0.6 inch=1.5cm

"R=\frac{5×10^{12}}{2π ×1} × In\frac{1.5}{1.27}"

"R=7.95\u00d710^{11}\u00d70.17=1.31\u00d710^{11}\u03a9\/m"